I was able to get the answer for this word problem but went the long way to get it. How would u work the problem? This is from an accelerated computer math program at my child's school & he doesn't have a book. What element of math is this called?
The question is: 8 teams are playing in a tournament. In the first round, each team plays each of the other teams once. How many games are played in the first round of the tournament?

Answer by richard1234(7193)

This would probably be classified in a field of math called "combinatorics."

Each team plays 7 games, so intuitively the number of games would be 8*7 = 56. However, we are counting each game twice (e.g. team 1 & team 6, team 6 & team 1) so we divide by 2 to get 28. Using combinatorics, the number of games would be 8C2 = 8*7/2! = 28.